# PTM-Psi on the Cloud

### v1

@@ 3.2 The Local PTM Perturbations Impact the Dynamics of Proximal Modules
Figure 4: (A) Comparison of residue-specific secondary structure preferences averaged over trajectories under various PTM perturbations. The descriptions of the PTM perturbation under the reference number (Ref 0-15 ) can be found in Table 1. Ref 0 is the native configuration. (B) Comparison of per-residue RMSF analysis under various PTM perturbations. We grouped the residue indices into GAP2, PRK, and CP12 modules with a chain ID in the same color schemes as those in Figure 1. We followed the alphabetical order of the chain IDs according to their chain labels in PDB ID: 6GVE. The intensity of RMSF is provided in the color bar. RMSF greater than 0.5nm is shown in deep blue.
Figure 4A show the averaged secondary structure propensity over the entire simulation under PTM perturbations. Compared to the native setting (Ref 0), the changes in the secondary structure motifs (e.g., $\alpha$ helix, $\beta$-sheet, coil/disordered) under various PTM perturbations are minimal, in agreement with the observation of the dark complex’s global dynamics in the previous section.
When we calculated the Residue-Wise Flexibility Profile (RMSF) to evaluate the structural fluctuations under PTM perturbations in 4B. The darker the blue color, the higher the RMSF, showing these residues are structurally flexible. This figure highlights the flexible or rigid regions of the dark complex under PTM perturbations. In the native configuration (Ref 0), Each PRK and CP12 has two pairs of disulfide bridges. Loops and termini often exhibit elevated mobility, while secondary structures such as helices and sheets remain relatively stable with low RMSF values.
Under PTM perturbations, we observed the emergence of narrow segments with high RMSF, not only in the regions near PTM perturbations, but also at the interface of proximal modules. For example, the RMSF of Ref 1, 2, and 3 (colored in gray) shows the impact of the perturbation of an oxidized cysteine from eight subunits that form two GAP2 tetrameters (Chain ID labeled with “ADEF” and “IKNO”). There are a few pronounced residual fluctuations within GAP2, as well as CP12 and PRK modules in the dark complex. When the two disulfide bridges in PRK (Chain ID labaled with “JG” and “BL”) are each perturbed with thiols, for Ref 4-9 (colored in magenta), a few prominent RMSF values emerge not only within PRK modules, but also near CP12 and GAP2. Similarly, when the two disulfide bridges in CP12 are each perturbed with thiols, for Ref 10-15 (colored in orange), high RMSF values appear only within CP12 modules, but also near GAP2 and PRK. In summary, while we computationally perturbed PTM one module at a time, we observed that its local impact propagates to the proximal modules of the whole dark complex.
@@ 3.2.1 The Structural Variability of Ensemble Conformations under PTM Perturbation is revealed by PCA
While single-variable metrics (e.g. Rg or RMSD) report isolated aspects of motion, they are unable to explain the relationships between multiple variables. By pooling all measurements into one feature matrix per frame, followed by PCA, we revealed the variability of PTM perturbation projected on the first principal component. We projected multiple features on the first principal component, allowing us to quantitatively rank the variability caused by PTM perturbation on the conformational ensemble.
[…]
@@ 3.3 Kinetic Transition Network
While principal component analysis (PCA) captures the major directions of conformational variance across molecular dynamics (MD) simulations, it has yet to resolve how or when a protein moves between those conformations over time. Therefore, we employed the KTN approach to analyze the pathways and transition states of molecular systems from the trajectories of MD by mapping the configurations onto the stable states or the transition pathways on an energy landscape. KTN not only enables us to track the change in dynamic accessibility and the interconversion between basins on the energy landscape under PTM perturbations, but also provides the explainable features that contribute to such perturbations.
@@
We begin by selecting the top-ranked reference in Table 1 from the PCA where it exhibits substantial deviation upon PTM perturbation along the first principal component (PC1). As an example for demonstration, we selected the PTM perturbation in CYS77 of GAP2 because it is not involved in a disulfide bridge as the cysteines in other two modules. We use Ref 3 because it has a higher variance than Ref 1 and Ref 2. It is expected because the size of SSG PTM of Ref 3 is the largest among the three thiol PTMs. We projected the feature descriptors of Ref 3 on the first and the second principal component, PC1 and PC2 ( accounting for $>90\%$ of cumulative variance). We applied the $k$-means clustering of $k=3$ for three metastable states. We then constructed a Markovian transition matrix $T$ from successive frame cluster assignments and computed stationary populations $\pi$, residence times $\tau$, and the mean first-passage times $\mathrm{MFPT}$ for each cluster. Please see Methods for the descriptions and definitions.
@@
Figure 6 shows the resultant kinetic transition network constructed from a 100 ns trajectory of the S-glutathionylated CYS77 dark complex (Ref 3), revealing three metastable basins. The largest basin (cluster ID“0”) captures 45% of the ensemble and exhibits a mean residence time, $\tau$, of approximately 0.24 ns. This cluster shows a “breathing” motion of the three GAP2, PRK, and CP12 modules, resulting in a slight increase in Rg and a decrease in compactness. The second most populated state (cluster ID“1”) accounts for 34% of the frames and has $\tau$ = 0.25 ns. It involves the local movement of PTMed CYS77 and nearby residues from both CP12 and GAP2 modules, increasing in $n_{PTM-residue}$. The third basin (cluster ID“2”) comprises 21% of the conformational ensemble and $\tau\approx 0.30$ ns. It involves helix-coil transitions in the CP12 module.
@@
These three clusters are similarly populated, and they are each attributed to local minima separated by a kinetic barrier on an energy landscape. The probability of passing over a kinetic barrier from one minimum to another, however, varies. The most frequent transitions occur between the two basins of ID 0 and ID 1, with mean first-passage times (MFPTs) of roughly 0.44 ns. The PTM perturbation on CYS77 of GAP2 quickly propagates into a global “breathing” motion within the three modules in the dark complex with a relatively high transition probability of p=0.29. In contrast, transitions from either basin ID 0 or ID 1 to basin ID 2 are relatively slow by a factor of  4, indicating a higher barrier for loop reorganization in CP12 than dynamics between module movements upon the perturbation of S-glutathionylation PTM of CYS77 on GAP2.
@@
Figure 6: The kinetic transition network was constructed for system Ref 3, where CYS77 of GAP2 was perturbed by SSG PTMs. The size of the node size is proportional to the stationary probability $\pi_{j}$. The width of an edge is proportional to the transition probability $T_{ij}$ (labels = $\mathrm{MFPT}_{i\to j}$ in ns). Behind each node is a representative structure near the centroid of a node. We highlighted the configurational differences in a certain region, including the loop of CP12 close to GAP2’s CYS77 and NAD-binding sites. For visual guidance, NAD is colored blue, CYS77 in yellow, disulfide bridges in PRK in cyan, and CP12 in green.
@@ 3.4 Thermodynamic Integrations
Figure 7: Relative binding free energy calculations (RBFE) from the TI feature in the PTM-psi package. This calculation involves transforming from CYS77 in chain A of GAP2 to SSG (Ref 3, A) and all CYS77 residues in chain A, D, E, F, I, K, N, and O of GAP2 to SSG (Ref 3, All). A positive RBFE value indicates that modifications from either single or all eight CYS77 residues to SSG are energetically unfavorable in GAP2.
One of the features of PTM-psi is its ability to perform relative binding free energy (RBFE) calculations using TI. We continued to focus on Ref 3 and calculated the relative binding free energy (RBFE) to assess the energetic impact of PTMs. Specifically, we performed RBFE calculations, transforming CYS77 in chain A of GAP2 to SSG individually, as well as transforming CYS77 across all eight chains of GAP2 tetramers to SSG. As illustrated in Figure 7, the RBFE calculations revealed that the PTM of a single CYS77 residue to SSG is energetically unfavorable by 16.38 kcal/mol, while the simultaneous modification of all eight CYS77 residues is even more unfavorable, with a total free energy change of 130.71 kcal/mol. It implies that SSG PTM at GAP2 might strategically destabilize the entire dark complex via propagated dynamics revealed by the kinetic transition network (KTN) model. The TI analysis offers thermodynamics insights into the impact of redox on the assembly of dark complex.
3.4.1 Software Optimizations vs Hardware Efficiencies for PTM-Psi on the Cloud
PTM-Psi relies heavily on GROMACS7 and NWChem6 for the most computationally demanding tasks in the workflows. Both software stacks have different capabilities in terms of the hardware requirements and scalability of the algorithms being used. As a consequence, adaptation of each part of the workflows to the appropriate hardware stack in the Cloud environment is needed. This adaptation has different objectives, but the most important ones are to minimize the time to obtain a given result while keeping costs to a minimum. For example, as of GROMACS version 2024.5, GPU offloading27, 28 of several tasks needed for free-energy perturbation was not supported.29 The end result is that free-energy perturbation runs requesting vGPUs can be considerably slower and much more expensive than those requesting only vCPUs.
In contrast, cloud computing is the on-demand delivery of computing power, databases, storage, applications, and other information technology resources via the internet with a pay-as-you-go pricing scheme. Large cloud service providers (CSP) like Microsoft Azure, Amazon Web Services (AWS), and Google Cloud Platform offer virtually limitless computing power distributed across their multi-geographic region data centers. Both High-Performance Computing (HPC) – where a large number of processors are needed to solve one particularly complex task – and High-Throughput Computing (HTC) – where a large number of simple and computationally independent tasks can be executed in a massively parallel fashion–can be tackled with the offerings from all three cloud computing providers. In the cloud, the hardware can be configured to perfectly suit the application requirements, allowing for the perfect match between hardware and workload.
@@ 4 Conclusions
We extended PTM-Psi on the cloud to make it fully scalable to a proteomic level and configurable for high-performance computing depending on the hardware. Together with the analysis tools, including PCA and KTN that provide insight into kinetic modeling from massive data, we can further advance scientific discovery by formulating an experimentally testable hypothesis of targeting site-specific cysteines that drive the assembly of the dark complex for redox signaling. Its far-reaching impact would be to control the yield of the Calvin-Benson Cycle pathway for sugar production with the approach of synthetic biology.
@@ 5 Data and Software Availability
The PTM-Psi package is made freely available under the terms of a modified 3-clause BSD license. The source code and installation instructions can be found in the GitHub repository https://github.com/pnnl/PTMPSI. The raw data, including molecular dynamics trajectories and topology files, will be made publicly available through PNNL’s DataHub repository.

### v2

@@ 3.4 The Local PTM Perturbations Impact the Dynamics of Proximal Modules
Figure 4: (A) Comparison of residue-specific secondary structure preferences averaged over trajectories under various PTM perturbations. The descriptions of the PTM perturbation for the simulation systems (Sim.0-15 ) can be found in Table 1. Sim.0 is the native configuration. (B) Comparison of per-residue RMSF analysis under various PTM perturbations. We grouped the residue indices into GAP2, PRK, and CP12 modules with a chain ID in the same color schemes as those in Figure 1. We followed the alphabetical order of the chain IDs according to their chain labels in PDB ID: 6GVE. The intensity of RMSF is provided in the color bar. RMSF greater than 0.5nm is shown in deep blue.
Figure 4A show the averaged secondary structure propensity over the entire simulation under PTM perturbations. Compared to the native setting (Sim.0), the changes in the secondary structure motifs (e.g., $\alpha$ helix, $\beta$-sheet, coil/disordered) under various PTM perturbations are minimal, in agreement with the observation of the dark complex’s global dynamics in the previous section.
When we calculated the RMSF to evaluate the residue-wise flexibility profile under PTM perturbations in 4B. The darker the blue color, the higher the RMSF, showing these residues are structurally flexible. This figure highlights the flexible or rigid regions of the dark complex under PTM perturbations. In the native configuration (Sim.0), Each PRK and CP12 has two pairs of disulfide bridges. Loops and termini often exhibit elevated mobility, while secondary structures such as helices and sheets remain relatively stable with low RMSF values.
Under PTM perturbations, we observed the emergence of narrow segments with high RMSF values, not only in the regions near PTM perturbations, but also at the interface of proximal modules. For example, the RMSF of Sim.1, 2, and 3 (colored in gray) shows the impact of the perturbation of an oxidized cysteine from eight subunits that form two GAP2 tetrameters (Chain ID labeled with “ADEF” and “IKNO”). There are a few pronounced residual fluctuations within GAP2, as well as CP12 and PRK modules in the dark complex. When the two disulfide bridges in PRK (Chain ID labaled with “JG” and “BL”) are each perturbed with thiols, for Sim.4-9 (colored in magenta), a few prominent RMSF values emerge not only within PRK modules, but also near CP12 and GAP2. Similarly, when the two disulfide bridges in CP12 are each perturbed with thiols, for Sim.10-15 (colored in orange), high RMSF values appear only within CP12 modules, but also near GAP2 and PRK. In summary, while we computationally perturbed PTM one module at a time, we observed that its local impact propagates to the proximal modules of the whole dark complex.
@@ 3.4.1 The Structural Variability of Ensemble Conformations under PTM Perturbation is revealed by PCA
While single-variable metrics (e.g., $R_{g}$ or RMSD) report isolated aspects of motion, they are unable to explain the relationships between multiple variables. By pooling all measurements into one feature matrix per frame, followed by PCA, we revealed the variability of PTM perturbation projected on the first principal component. We projected multiple features on the first principal component, allowing us to quantitatively rank the variability caused by PTM perturbation on the conformational ensemble.
As detailed in Methods, we performed a global PCA over all 16 simulations and computed the $\langle\mathrm{PC1}\rangle$. The first principal component PC1 represents the direction of maximal variance in the data. It is the axis where the data spreads out the most. To emphasize the variability compared with the native state, using Sim. 0 as a baseline, we plot the $\Delta\langle\mathrm{PC1}\rangle$ score in Figure 5. It shows that the PTM perturbation drives the movement of GAP2 and CP12 opposite to the baseline, while that of PRK in the same direction as the baseline along PC1. In other words, although PTM perturbations have not altered local structures in each module significantly, they have indeed changed the global dynamics of the dark complex.
Figure 5: (A) $\Delta\langle\mathrm{PC1}\rangle$ Score for each simulation system perturbed by PTMed (Sim. 0 is the baseline). Positive (negative) values show positive (negative) correlation with respect to the Sim. 0. (B) A structural representation of GAPDH/CP12/PRK megastructure, with SSG modification in GAP2 modules of PDB: 6GVE (Sim. 3) in Table 1. PTM perturbations in the GAP2 modules appear in gray, PRK in pink, CP12 in magenta, oxidized cysteine (CYS78) residue of GAP2 in yellow, NAD cofactor in blue, cysteine residues forming disulfide bridges in cyan in PRK and green in CP12, respectively.
@@ 3.5 Kinetic Transition Network
While PCA captures the major directions of conformational variance across MD simulations, it has yet to resolve how or when a protein moves between those conformations over time. Therefore, we employed the KTN approach to analyze the pathways and transition states of molecular systems from the trajectories of MD by mapping the configurations onto the stable states or the transition pathways on an energy landscape. KTN not only enables us to track the change in dynamic accessibility and the interconversion between basins on the energy landscape under PTM perturbations, but also provides the explainable features that contribute to such perturbations.
@@
We begin by selecting the top-ranked simulation system in Table 1 from the PCA where it exhibits substantial deviation upon PTM perturbation along the PC1. As an example for demonstration, we selected the PTM perturbation in CYS77 of GAP2 because it is not involved in a disulfide bridge as the cysteines in other two modules. We use Sim.3 because it has a higher variance than Sim.1 and Sim.2. It is expected because the size of SSG PTM of Sim.3 is the largest among the three thiol PTMs. We projected the feature descriptors of Sim.3 on the first and the second principal components, PC1 and PC2 ( accounting for $90\%$ of cumulative variance). We applied the $k$-means clustering of $k=3$ for three metastable states. We then constructed a Markovian transition matrix $T$ from successive frame cluster assignments and computed stationary populations $\pi$, residence times $\tau$, and the mean first-passage times $\mathrm{MFPT}$ for each cluster. Please see Methods for the descriptions and definitions.
@@
Figure 6 shows the resultant KTN constructed from a 100 ns trajectory of the S-glutathionylated CYS77 dark complex (Sim.3), revealing three metastable basins. The largest basin (cluster ID“0”) captures 45% of the ensemble and exhibits a mean residence time, $\tau$, of approximately 0.24 ns. This cluster shows a “breathing” motion of the three GAP2, PRK, and CP12 modules, resulting in a slight increase in $R_{g}$ and a decrease in compactness. The second most populated state (cluster ID“1”) accounts for 34% of the frames and has $\tau$ = 0.25 ns. It involves the local movement of PTMed CYS77 and nearby residues from both CP12 and GAP2 modules, increasing in $n_{PTM-residue}$. The third basin (cluster ID“2”) comprises 21% of the conformational ensemble and $\tau\approx 0.30$ ns. It involves helix-coil transitions in the CP12 module.
@@
These three clusters are similarly populated, and they are each attributed to local minima separated by a kinetic barrier on an energy landscape. The probability of passing over a kinetic barrier from one minimum to another, however, varies. The most frequent transitions occur between the two basins of ID 0 and ID 1, with MFPTs of roughly 0.44 ns. The PTM perturbation on CYS77 of GAP2 quickly propagates into a global “breathing” motion within the three modules in the dark complex with a relatively high transition probability of p=0.29. In contrast, transitions from either basin ID 0 or ID 1 to basin ID 2 are relatively slow by a factor of  4, indicating a higher barrier for loop reorganization in CP12 than dynamics between module movements upon the perturbation of S-glutathionylation PTM of CYS77 on GAP2.
@@
Figure 6: The KTN was constructed for system Sim.3, where CYS77 of GAP2 was perturbed by SSG PTMs. The size of the node size is proportional to the stationary probability $\pi_{j}$. The width of an edge is proportional to the transition probability $T_{ij}$ (labels = $\mathrm{MFPT}_{i\to j}$ in ns). Behind each node is a representative structure near the centroid of a node. We highlighted the configurational differences in a certain region, including the loop of CP12 close to GAP2’s CYS77 and NAD-binding sites. For visual guidance, NAD is colored blue, CYS77 in yellow, disulfide bridges in PRK in cyan, and CP12 in green.
@@ 3.6 Thermodynamic Integrations
Figure 7: The RBFE calculations from the TI feature in the PTM-Psi package. This calculation involves transforming from CYS77 in chain A of GAP2 to SSG (Sim.3, A) and all CYS77 residues in chain A, D, E, F, I, K, N, and O of GAP2 to SSG (Sim.3, All). A positive RBFE value indicates that modifications from either single or all eight CYS77 residues to SSG are energetically unfavorable in GAP2.
One of the features of PTM-Psi is its ability to perform RBFE calculations using TI. We continued to focus on Sim.3 and calculated the RBFE to assess the energetic impact of PTMs. Specifically, we performed RBFE calculations, transforming CYS77 in chain A of GAP2 to SSG individually, as well as transforming CYS77 across all eight chains of GAP2 tetramers to SSG. As illustrated in Figure 7, the RBFE calculations revealed that the PTM of a single CYS77 residue to SSG is energetically unfavorable by 16.38 kcal/mol, while the simultaneous modification of all eight CYS77 residues is even more unfavorable, with a total free energy change of 130.71 kcal/mol. It implies that SSG PTM at GAP2 might strategically destabilize the entire dark complex via propagated dynamics revealed by the KTN model. The TI analysis offers thermodynamics insights into the impact of redox on the assembly of dark complex.
@@ 4 Conclusions
We extended PTM-Psi on the cloud to make it fully scalable to a proteomic level and configurable for HPC depending on the hardware. Together with the analysis tools, including PCA and KTN that provide insight into kinetic modeling from massive data, we can further advance scientific discovery by formulating an experimentally testable hypothesis of targeting site-specific cysteines that drive the assembly of the dark complex for redox signaling. Its far-reaching impact would be to control the yield of the CBC pathway for sugar production with the approach of synthetic biology.
5 Conflict of Interest Statement
The authors declare no conflicts of interest regarding this manuscript.
6 Author Contribution
M.S.C and D.M.-R. designed the research. P.R. and A.A. contributed to code adaptation. D.M.-R. and H. K. executed the simulations. S.S., H.K., and M.L. analyzed the data. All authors contributed to writing the manuscript.
@@ 7 Data and Software Availability
The PTM-Psi package is made freely available under the terms of a modified 3-clause BSD license. The source code and installation instructions can be found in the GitHub repository https://github.com/pnnl/PTMPSI. The raw data, including molecular dynamics (MD) trajectories and topology files, will be made publicly available through Pacific Northwest National Laboratory’s DataHub repository.